# Stochastics of shapes and Kunita flows

### v1

@@ Abstract.
Media Summary
Biology has many shapes - from the curves of butterfly wings to the complex forms of human organs. These shapes can change over time through evolution, but describing and analyzing how shapes change randomly is an intricate mathematical problem.
In this paper, we focus on mathematical tools to properly describe how shapes can change randomly over time. We show how Kunita flows can be used to model shape changes in a way that respects the natural properties of shapes. This approach allows us to track and measure shape variations in a consistent way, regardless of how we choose to represent the shapes mathematically.
Our work provides a foundation for studying shape evolution in fields like biology, where understanding how animal and plant forms change through random evolutionary processes is crucial. We also show how these mathematical tools can be combined with modern statistical methods to incorporate the information in real shape data.
@@ 1. Introduction
Statistics of shapes has been treated extensively in the literature, both because of the mathematical intricacies resulting from working with shapes as inherently non-linear objects, and because of the many applications in e.g. biology and medical imaging. Shape spaces span from landmark based representation by [kendallShapeManifoldsProcrustean1984] over elastic metrics ([mioShapePlaneElastic2007, younesMetricShapeSpace2008]) to infinite-dimensional diffeomorphism based representations ([younesShapesDiffeomorphisms2010]) with many additional models and shape spaces in between. This span includes many perspectives on how shapes should be modeled, how metrics on shape spaces can be constructed, and how statistics of shapes can be performed.
@@ 2.1. Outer shape models
for covectors $p,\tilde{p}\in(\mathbb{R}^{d})^{n}$ and a Greens kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ of $L$. Geodesic equations for such landmark equations can be found by defining the Hamiltonian $H(x,p)=\frac{1}{2}\left<p,p\right>_{x}$ using the cometric and solving the Hamilton-Jacobi equation
@@ 3. An axiomatic approach to shape stochastics
Below, we discuss and formalize these properties. See also Figure 2 for a schematic illustration of the intuition behind the properties. We assume the processes are defined on a standard probability space $(\Omega,\mathcal{F},\mathbb{P})$. As we in the end are interested in the statistical properties of the shape processes, the convergence properties are formalized as statements in distribution. When we have a map $f:\mathcal{S}\to\mathcal{S}_{1}$ between two shape spaces, we will use the notation $f_{*}s_{t}$ for the pushforward of a process $s_{t}$ in $\mathcal{S}$ by $f$. For each $t$, this means that the random variable $s_{t}$ is mapped to the random variable $f\circ s_{t}$, i.e. we just apply $f$ to the shape process. The law of $s_{1,t}$ is then $\mathbb{P}(f_{*}s_{t}\in A)=\mathbb{P}(s_{t}\in f^{-1}(A))$ for measurable subsets $A\subset\mathcal{S}_{1}$.
@@ 3.1. Representation and discretization independence
A common approach of defining shape dynamics is to define the processes directly on the representation space, for example on landmark configurations by perturbing each landmark. However, this often leads to a dependence on the representation, e.g. the specific number of landmarks. Whether a surface is represented as a surface, by a set of landmarks on the surface, or as an image from which the surface is segmented, the dynamics should ideally be equivalent up to the loss of information specific to each representation. The property covers this property.
@@ 3.3. Rigid transformation and reparametrization equivariance
Let $\operatorname{SE}(d)$ be the special Euclidean group of translations and rotations acting on the shape space $\mathcal{S}$ with action $g.s$, $g\in\operatorname{SE}(d)$. Let $s_{t}^{s_{0}}$ be a stochastic process in $\mathcal{S}$ started at $s_{0}\in\mathcal{S}$. We say that the process is equivariant to the action of $\operatorname{SE}(d)$ if the law of $s_{t}^{g.s_{0}}$ is the same as the law of $g.s_{t}^{s_{0}}$ for $s_{0}\in\mathcal{S}$ and $g\in\operatorname{SE}(d)$.
@@ 4. Kunita flows
We now focus on a concrete example of a family of processes satisfying the properties (1)-(4), the Kunita flows.
@@
Under additional technical assumptions, $\phi_{s,t}$ can be shown to be a Brownian flow of homeomorphisms meaning that 1) $\phi_{s,t}(x)$ is continuous in $s,t,x$ a.s.; and 2) $\phi_{s,t}$ is a homeomorphism a.s. If $\phi_{s,t}$ is a $C^{l}$, $l\in\mathbb{N}$, diffeomorphism a.s., then $\phi_{s,t}$ is called a Brownian flow of $C^{l}$ diffeomorphisms.
@@
Let $\phi_{s,t}$ be a Brownian flow of homeomorphisms. If $b_{t}$ and $a_{t}$ are $C^{l}$ in $x$ and $y$ with bounded $l$th order derivatives in both $x$ and $y$ simultaneously, then $\phi_{s,t}$ is a Brownian flow of $C^{l-1}$ diffeomorphisms.
@@ 4.1. Semimartingales and stochastic integrals
Let $\alpha,\beta$ be such that
$V(x,t)$ $=\int_{0}^{t}\beta(x,r)dr,$
$\left<M(x,t),M(y,t)^{*}\right$> $=\int_{0}^{t}\alpha(x,y,r)dr\ .$
Then $\alpha,\beta$ are called the local characteristics of the semimartingale.
@@
If the local characteristics $\alpha,\beta$ are Lipschitz continuous with linear growth, then (2) has a unique solution with a modification that is a stochastic flow of diffeomorphisms.
@@ 4.2. Hilbert space formulation
$\beta(x,t)$ $=0$
$\alpha(x,y,t)$ $=g(x,y).$
This formulation gives a direct correspondence between the kernel $k$, its square $g$ and the covariance characteristic $\alpha$ of the flow. The kernel $k$ can be chosen to satisfy the required smoothness conditions to ensure that the resulting flow is a flow of diffeomorphisms. As $k$ is often translation and rotation invariant and on scalar form, one can write $k(x,y)=k(r)\operatorname{Id}_{d}$ where $r=\|x-y\|$ and $k$ is now interpreted as a function $k:\mathbb{R}\to\mathbb{R}$.
@@
which results in $g_{\text{SE}}(r)=\alpha^{2}\pi^{d/2}\sigma^{d}e^{-\frac{r^{2}}{4\sigma^{2}}}$ being on the squared exponential form as well. Another common choice are Green’s functions of Laplacian and higher order Sobolev operators. For such an operator $L=\alpha^{-1}(\operatorname{Id}-\sigma^{2}\Delta)^{c}$, we have
@@
The Hilbert space formulation in addition gives a natural way to parametrize the flow applied to curves using other bases, e.g. Fourier space representations. This is developed in [bakerConditioningNonlinearInfinitedimensional2024]. With a Fourier basis $\{e_{n}\}_{n=1}^{\infty}$, $e_{n}(x)=e^{inx}$ and $g_{l,m}(x_{1},x_{2})=e^{ilx_{1}+imx_{2}}$, writing the SDE (3) into basis elements gives
@@ 4.4. Kunita flows and shape stochastics
Let $\phi_{t}$ be a Kunita flow with local characteristics $\alpha,\beta$ such that $\phi_{t}$ is a flow of $C^{l}(\mathbb{R}^{d})$-diffeomorphisms. Let $\phi\in C^{l}(\mathbb{R}^{d})$ act on the shape space $\mathcal{S}$ by $\phi.s_{0}$, $s_{0}\in\mathcal{S}$. Then $\phi_{t}$ induces a shape process $s_{t}=\phi_{t}.s_{0}$ satisfying the properties (1) and (2). If the local characteristics are invariant to rigid motions, then property (3) holds.
@@
We now turn to discretization consistency. The law of a Kunita flow is determined by its finite dimensional distributions. This is equivalent to its motion being determined by the motion of any finite set of landmarks, i.e. that its action $\phi_{t}.s$ is known for all $s\in\mathcal{S}_{n}$ for the landmark configuration space $\mathcal{S}_{n}$ and all $n$. However, since the flow is characterized by its local characteristics, the law of the Kunita flow is actually determined only by its 2-point motions, i.e. if its action on $\mathcal{S}_{2}$ is known, see [kunitaStochasticFlowsStochastic1997, setion 4.2]. Referring to the discretization property (4), we see that the law of $\phi$ can be recovered if the family of finite dimensional discretizations $\mathcal{P}$ includes all 2-point pairs on $\mathbb{R}^{d}$.
@@ 4.5. Shape variance
4.6. Ito vs. Stratonovich
We have described Kunita flows on Ito form, but they can equivalently be formulated in Stratonovich form with Stratonovich integral $\circ dW_{t}$ and with changed drift originating from the Ito-Stratonovich correction term. Because mappings of Stratonovich processes obey the chain rule, the Stratonovich formulation is sometimes preferable. One example is the stochastic EPDiff shape flows and the fluid dynamics processes from which they arise ([holmVariationalPrinciplesStochastic2015]), where the Stratonovich version is used, see section 5. Shape flows induced by an Ito Kunita flow may receive an additional drift term involving the derivative of the action when written as an SDE directly on the shape space, in contrast to Stratonovich flows that, because of the chain rule, will not receive the additional drift. On the other hand, the Ito formulation can be easier to interpret and implement.
@@ 6.2. Conditioning in infinite dimensions
To move from landmarks to continuous shapes, we need to extend the finite dimensional constructions above to infinite dimensions. In the linear case, this has been treated in [goldysOrnsteinUhlenbeckBridge2008, dinunnoSPDEBridgesObservation2023]. However, as we have seen, shape stochastic processes generally depend non-linearly on the states (property (2) in section 3). For strong solutions of (linear and non-linear) Hilbert space-valued SDEs, [bakerConditioningNonlinearInfinitedimensional2024] develop a Doob’s $h$-transform for conditioning processes to end in a given subset. Hence viewing the SDE via the Hilbert space formulation, we can use this to construct a conditioned process.
@@ 6.3. Inference
Emulating a phylogenetic context where the shapes exhibit random changes through evolution and where species branches over time, we assume a root butterfly wing shape is affected by a stochastic process that branches according to a simple phylogenetic tree. The stochastic process is induced by a Kunita flow with Matérn kernel. When sampling the forward process, we get observations of the shapes at the leafs of the tree. Denoting such a set of leaf observations by $s_{1},\ldots,s_{N}$, we can use the guided bridge scheme to estimate the parameters of the process. This is done by using the guiding process (9) along each edge of the tree. This forward sampling is interchanged with a backwards filtering step that, for a given sample of parameters, propagates from leaves upwards in the tree the information necessary for guiding the bridge, i.e. the parameters of the auxiliary process $\tilde{p}_{t\to T}(x,y)$ in (9). The backward filtering follows the backward filtering scheme of [meulenBackwardFilteringForward2025]. These steps are repeated in an MCMC loop that iteratively proposes new values for the parameters $\alpha$ and $\sigma$ in between the forwards sampling and backwards filtering steps.

### v2

@@ 1. Introduction
Statistics of shapes has been treated extensively in the literature, both because of the mathematical intricacies resulting from working with shapes as inherently non-linear objects, and because of the many applications in e.g. biology and medical imaging. Shape spaces span from landmark based representation by Kendall ([kendallShapeManifoldsProcrustean1984]) over elastic metrics ([mioShapePlaneElastic2007, younesMetricShapeSpace2008]) to infinite-dimensional diffeomorphism based representations ([younesShapesDiffeomorphisms2010]) with many additional models and shape spaces in between. This span includes many perspectives on how shapes should be modeled, how metrics on shape spaces can be constructed, and how statistics of shapes can be performed.
@@ 2.1. Outer shape models
for covectors $p,\tilde{p}\in(\mathbb{R}^{d})^{n}$ and a Green’s kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ of $L$. Geodesic equations for such landmark equations can be found by defining the Hamiltonian $H(x,p)=\frac{1}{2}\left<p,p\right>_{x}$ using the cometric and solving the Hamilton-Jacobi equation
@@ 3. An axiomatic approach to shape stochastics
Below, we discuss and formalize these properties. See also Figure 2 for a schematic illustration of the intuition behind the properties. We assume the processes are defined on a standard probability space $(\Omega,\mathcal{F},\mathbb{P})$. As we in the end are interested in the statistical properties of the shape processes, the convergence properties are formalized as statements in distribution. When we have a map $f:\mathcal{S}\to\mathcal{S}_{1}$ between two shape spaces, we will use the notation $f_{*}s_{t}$ for the pushforward of a process $s_{t}$ in $\mathcal{S}$ by $f$. For each $t$, this means that the random variable $s_{t}$ is mapped to the random variable $f\circ s_{t}$, i.e. we just apply $f$ to the shape process. The law of $f_{*}s_{t}$ is then $\mathbb{P}(f_{*}s_{t}\in A)=\mathbb{P}(s_{t}\in f^{-1}(A))$ for measurable subsets $A\subset\mathcal{S}_{1}$.
@@ 3.1. Representation and discretization independence
A common approach to define shape dynamics is to define the processes directly on the representation space, for example on landmark configurations by perturbing each landmark. However, this often leads to a dependence on the representation, e.g. the specific number of landmarks. Whether a surface is represented as a surface, by a set of landmarks on the surface, or as an image from which the surface is segmented, the dynamics should ideally be equivalent up to the loss of information specific to each representation. The first property covers this representation independence.
@@ 3.3. Rigid transformation and reparametrization equivariance
Let $\operatorname{SE}(d)$ be the special Euclidean group of translations and rotations acting on the shape space $\mathcal{S}$ with action $g.s$, $g\in\operatorname{SE}(d)$. Let $s_{t}^{s_{0}}$ be a stochastic process in $\mathcal{S}$ started at $s_{0}\in\mathcal{S}$. The process is equivariant to the action of $\operatorname{SE}(d)$ on $s_{0}$ if the law of $s_{t}^{g.s_{0}}$ is the same as the law of $g.s_{t}^{s_{0}}$ for $g\in\operatorname{SE}(d)$.
@@ 4. Kunita flows
We now focus on a concrete example of a family of processes satisfying the properties (1)-(4), shape processes induced by Kunita flows.
@@
Under additional technical assumptions, $\phi_{s,t}$ can be shown to be a Brownian flow of homeomorphisms meaning that 1) $\phi_{s,t}(x)$ is continuous in $s,t,x$ a.s.; and 2) $\phi_{s,t}$ is a homeomorphism a.s. If $\phi_{s,t}$ is a $C^{l}$, $l\in\mathbb{N}$, diffeomorphism a.s., then $\phi_{s,t}$ is called a Brownian flow of $C^{l}$-diffeomorphisms.
@@
Let $\phi_{s,t}$ be a Brownian flow of homeomorphisms. If $b_{t}$ and $a_{t}$ are $C^{l}$ in $x$ and $y$ with bounded $l$th order derivatives in both $x$ and $y$ simultaneously, then $\phi_{s,t}$ is a Brownian flow of $C^{l-1}$-diffeomorphisms.
@@ 4.1. Semimartingales and stochastic integrals
Let $a_{t},b_{t}$ be such that
$V(x,t)$ $=\int_{0}^{t}b_{r}(x)dr,$
$\left<M(x,t),M(y,t)^{*}\right$> $=\int_{0}^{t}a_{r}(x,y)dr\ .$
Then $a_{t},b_{t}$ are called the local characteristics of the semimartingale.
@@
If the local characteristics $a_{t},b_{t}$ are Lipschitz continuous with linear growth, then (2) has a unique solution with a modification that is a stochastic flow of diffeomorphisms.
@@ 4.2. Hilbert space formulation
$b_{t}(x)$ $=0$
$a_{t}(x,y)$ $=g(x,y).$
This formulation gives a direct correspondence between the kernel $k$, its square $g$ and the covariance characteristic $a_{t}$ of the flow. The kernel $k$ can be chosen to satisfy the required smoothness conditions to ensure that the resulting flow is a flow of diffeomorphisms. As $k$ is often translation and rotation invariant and scalar valued, one can write $k(x,y)=k(r)\operatorname{Id}_{d}$ where $r=\|x-y\|$ and $k$ is now interpreted as a function $k:\mathbb{R}\to\mathbb{R}$.
@@
which results in $g_{\text{SE}}(r)=\alpha^{2}\pi^{d/2}\sigma^{d}e^{-\frac{r^{2}}{4\sigma^{2}}}$ being a squared exponential as well. Another common choice are Green’s functions of Laplacian and higher order Sobolev operators. For such an operator $L=\alpha^{-1}(\operatorname{Id}-\sigma^{2}\Delta)^{c}$, we have
@@
The Hilbert space formulation in addition gives a natural way to parametrize the flow applied to curves using other bases, e.g. Fourier space representations. This is developed in [bakerConditioningNonlinearInfinitedimensional2024]. In $\mathbb{R}^{2}$ with a Fourier basis $\{e_{n}\}_{n=1}^{\infty}$, $e_{n}(x)=e^{inx}$ and $g_{l,m}(x_{1},x_{2})=e^{ilx_{1}+imx_{2}}$, writing the SDE (3) into basis elements gives
@@ 4.4. Kunita flows and shape stochastics
Let $\phi_{t}$ be a Kunita flow with local characteristics $a_{t},b_{t}$ such that $\phi_{t}$ is a flow of $C^{l}(\mathbb{R}^{d})$-diffeomorphisms. Let $\phi\in C^{l}(\mathbb{R}^{d})$ act on the shape space $\mathcal{S}$ by $\phi.s_{0}$, $s_{0}\in\mathcal{S}$. Then $\phi_{t}$ induces a shape process $s_{t}=\phi_{t}.s_{0}$ satisfying the properties (1) and (2). If the local characteristics are invariant to rigid motions, then property (3) holds.
@@
We now turn to discretization consistency. The law of a Kunita flow is determined by its finite dimensional distributions. This is equivalent to its motion being determined by the motion of any finite set of landmarks, i.e. that its action $\phi_{t}.s$ is known for all $s\in\mathcal{S}_{n}$ for the landmark configuration space $\mathcal{S}_{n}$ and all $n$. However, since the flow is characterized by its local characteristics, the law of the Kunita flow is actually determined only by its 2-point motions, i.e. if its action on $\mathcal{S}_{2}$ is known, see [kunitaStochasticFlowsStochastic1997, section 4.2]. Referring to the discretization property (4), we see that the law of $\phi$ can be recovered if the family of finite dimensional discretizations $\mathcal{P}$ includes all 2-point pairs on $\mathbb{R}^{d}$.
@@ 4.5. Shape variance
4.6. Itô vs. Stratonovich
We have described Kunita flows on Itô form, but they can equivalently be formulated in Stratonovich form with Stratonovich integral $\circ dW_{t}$ and with changed drift originating from the Itô-Stratonovich correction term. Because mappings of Stratonovich processes obey the chain rule, the Stratonovich formulation is sometimes preferable. One example is the stochastic EPDiff shape flows and the fluid dynamics processes from which they arise ([holmVariationalPrinciplesStochastic2015]), where the Stratonovich version is used, see section 5. Shape flows induced by an Itô Kunita flow may receive an additional drift term involving the derivative of the action when written as an SDE directly on the shape space, in contrast to Stratonovich flows that, because of the chain rule, will not receive the additional drift. On the other hand, the Itô formulation can be easier to interpret and implement.
@@ 6.2. Conditioning in infinite dimensions
To move from landmarks to continuous shapes, we need to extend the finite dimensional constructions above to infinite dimensions. In the linear case, this has been treated in [goldysOrnsteinUhlenbeckBridge2008, dinunnoSPDEBridgesObservation2023]. However, as we have seen, shape stochastic processes generally depend non-linearly on the states (property (2) in section 3). For strong solutions of (linear and non-linear) Hilbert space-valued SDEs, [bakerConditioningNonlinearInfinitedimensional2024] developed a Doob’s $h$-transform for conditioning processes to end in a given subset. Hence viewing the SDE via the Hilbert space formulation, we can use this to construct a conditioned process.
@@ 6.3. Inference
Emulating a phylogenetic context where the shapes exhibit random changes through evolution and where species branches over time, we assume a root butterfly wing shape is affected by a stochastic process that branches according to a simple phylogenetic tree. The stochastic process is induced by a Kunita flow with Matérn kernel. When sampling the forward process, we get observations of the shapes at the leaves of the tree. Denoting such a set of leaf observations by $s_{1},\ldots,s_{N}$, we can use the guided bridge scheme to estimate the parameters of the process. This is done by using the guiding process (9) along each edge of the tree. This forward sampling is interchanged with a backwards filtering step that, for a given sample of parameters, propagates from leaves upwards in the tree the information necessary for guiding the bridge, i.e. the parameters of the auxiliary process $\tilde{p}_{t\to T}(x,y)$ in (9). The backward filtering follows the backward filtering scheme of [meulenBackwardFilteringForward2025]. These steps are repeated in an MCMC loop that iteratively proposes new values for the parameters $\alpha$ and $\sigma$ of the Matérn kernel in between the forwards sampling and backwards filtering steps.